### v2

# Survey of Computerized Adaptive Testing: A Machine Learning Perspective

Abstract.
Computerized Adaptive Testing (CAT) provides an efficient and tailored method for assessing the proficiency of examinees, by dynamically adjusting test questions based on their performance. Widely adopted across diverse fields like education, healthcare, sports, and sociology, CAT has revolutionized testing practices. While traditional methods rely on psychometrics and statistics, the increasing complexity of large-scale testing has spurred the integration of machine learning techniques. This paper aims to provide a machine learning-focused survey on CAT, presenting a fresh perspective on this adaptive testing method. By examining the test question selection algorithm at the heart of CAT’s adaptivity, we shed light on its functionality. Furthermore, we delve into cognitive diagnosis models, question bank construction, and test control within CAT, exploring how machine learning can optimize these components. Through an analysis of current methods, strengths, limitations, and challenges, we strive to develop robust, fair, and efficient CAT systems. By bridging psychometric-driven CAT research with machine learning, this survey advocates for a more inclusive and interdisciplinary approach to the future of adaptive testing.
Computerized Adaptive Testing, Machine Learning, Cognitive Diagnosis, Intelligent Measurement, Question Selection
ccs: Applied computing Learning management systemsccs: Applied computing Computer-managed instruction
1. Introduction
Computerized Adaptive Testing (CAT), originating from psychometrics, allows for efficient assessment through personalized question selection (Chen et al., 2015a; Vie et al., 2017). It has been widely applied in high-stakes testing scenarios, including the Duolingo Test, GRE, TOEFL, and GMAT. The term “computerized” underscores its origins, reflecting a combination of computer science and intelligent measurement. As shown in Figure 1, in contrast to the one-for-all approach of traditional paper-and-pencil tests, CAT tailors the selection of questions to each examinee’s level of proficiency, thereby maximizing the accuracy of the assessment while minimizing the test length. In essence, CAT strives to answer a pivotal question about accuracy and efficiency: How to accurately estimate an examinee’s true proficiency by providing the minimal number of questions?
Figure 1. If (a) traditional paper-and-pencil test is “one-for-all”, then (b) CAT is “one-for-each”. Each examinee gets a personalized test that adapts to his/her proficiency level and knowledge, ensuring each question accurately assesses and challenges the examinee
CAT is a dynamic and interactive process between an examinee and a testing system. The CAT system includes four main components that take turns: At each test step, the Cognitive Diagnosis Model, as the user model, first uses the examinee’s previous responses to estimate their current proficiency, based on cognitive science or psychometrics (Ackerman et al., 2003). Then, the Selection Algorithm picks the next question from the Question Bank according to certain criteria (Lord, 2012; Chang and Ying, 1996; Bi et al., 2020). Most traditional criteria are statistical informativeness metrics, e.g., selecting the question whose difficulty matches the examinee’s current proficiency estimate, meaning the examinee has roughly a 50% chance of getting it right. The above process repeats until a predefined stopping rule is met. Throughout the assessment, Test Control governs various factors such as content balance, fairness, and robustness of the testing process. At the conclusion of the CAT, the final proficiency estimate—or the diagnostic report—is provided to the examinee as the outcome of this assessment.
Obviously, CAT represents a complex fusion of machine intelligence and assessment techniques. It needs to manage large question banks, adapt to varying examinee proficiencies, and real-time decision-making. Moreover, practical CAT also involves ensuring reliability, fairness, search efficiency, etc. These challenges make CAT a multifaceted decision-making problem that requires sophisticated solutions. Particularly, with the emergence of large-scale and diverse online testing platforms, traditional statistical methods often struggle to optimally select questions while addressing these multifarious factors. Machine learning (Jordan and Mitchell, 2015), with its ability to learn from data and make precise predictions, offers a viable solution to enhance test efficiency and accuracy within this volatile environment. Previous CAT surveys (Chang, 2015; Vie et al., 2017; Cheng, 2008; Mujtaba and Mahapatra, 2020) have largely revolved around the domains of statistics and psychometrics. Given CAT’s interdisciplinary nature, this paper seeks to understand and review methodologies from a machine-learning perspective, which is more approachable to a wider readership and sheds light on building strong CAT systems.
In the world of machine learning, CAT can be viewed as a parameter estimation problem with a focus on data efficiency (Mirzasoleiman et al., 2020; Zhuang et al., 2023): the goal is to sequentially select the smallest number of data samples (questions) from a pool of unlabeled data (the question bank) so that, after their labels (correct/wrong responses) are observed, the model’s latent parameters (examinee’s true proficiency $\theta_{0}$) can be estimated accurately. In recent years, there has been a growing interest in applying machine learning techniques to investigate the four components in CAT. For example, deep learning techniques diagnose examinee’s proficiency (Wang et al., 2020c) and automate question bank construction (Qiu et al., 2019); data-driven approaches optimize selection algorithms by learning from large-scale response data (Zhuang et al., 2022b; Li, 2020; Ghosh and Lan, 2021); machine learning enhances CAT’s reliability by addressing various factors (e.g., robustness, search efficiency) (Hong et al., 2023). Despite these efforts, a comprehensive survey that captures the breadth of CAT solutions from a machine-learning perspective is still lacking. Furthermore, the ongoing evolution of machine learning presents new aspects for CAT.
This paper will help identify the strengths and limitations of machine learning approaches in CAT, highlight the challenges, and provide guidance for future research directions. Given the fact that the selection algorithm is the core component for achieving the adaptivity, this survey mainly focuses on its advancements. The contributions of this paper are as follows:
  • To our knowledge, this represents the first attempt to comprehensively review CAT solutions through the lens of machine learning. By exploring the existing work in Cognitive Diagnosis Model, Selection Algorithms, Question Bank Construction, and Test Control, the paper offers a unified framework and encompasses the entire life cycle of the CAT system.
  • We summarize existing works and draw conclusions on the success and failure attempts of machine learning. We also discuss the crucial factors in building a reliable CAT, e.g., content balance, fairness, robustness, and search efficiency, often overlooked in prior reviews.
  • We have open-sourced extensible and unified implementations of existing CAT models and relevant resources at https://github.com/bigdata-ustc/EduCAT. This library aims to assist researchers in swiftly developing a CAT system, encouraging collaboration, and ultimately leading to more sophisticated and effective CAT systems.
The paper is organized as follows. In Section 2, we introduce the history and the contemporary trend of CAT. Then in Section 3, we provide the formulation of CAT’s task. After that, Section 4, 5, and 6 respectively review the existing methods for the cognitive diagnosis model, selection algorithms, and question bank construction. In Section 7, we summarize the key factors in the application of CAT. Section 8 discusses how to evaluate the CAT. We discuss the future directions in Section 9 and Section 10 concludes the paper.
2. Background of CAT
Evolution of CAT
The evolution of CAT is a fascinating journey through time, marked by significant milestones that have shaped the landscape of educational measurement.
  • The Rise of CAT. The story begins in the early 20th century with Alfred Binet, a psychologist who developed the first form of adaptive testing in 1905 as an intelligence test (Wainer et al., 2000). In the 1950s, computers began transforming adaptive testing into what we now call CAT. Groundbreaking work throughout the 1970s and 1980s solidified the foundation of CAT, with subsequent refinements enhancing precision and efficiency (Sands et al., 1997). A pivotal advancement was the integration of the Rasch model from Item Response Theory (IRT) (Roskam and Jansen, 1984), which allowed CAT to adaptively match question difficulty to a test-taker’s estimated proficiency level, thereby improving the accuracy of proficiency assessments.
  • Widespread Adoption. In the 1990s, the advent of the internet made CAT more accessible, leading to its adoption by major tests like the GRE and TOEFL. While these tests have evolved from the classic CAT paradigm, adaptive testing principles continue to inform their current testing models. Various statistical methods—Bayesian, maximum likelihood, or maximum information methods–are employed to optimize the testing experience. During this time, CAT emerged as a dominant topic in educational measurement forums. For example, at the 1999 annual meeting of the National Council on Measurement in Education, a quarter of the papers presented were devoted to CAT (Meijer and Nering, 1999).
  • CAT Meets Machine Learning. Current CAT research spans a wide range of topics, including the development of question banks, question selection, proficiency estimation, and various issues related to test security and reliability. They are critical to ensuring that CAT remains a reliable, valid, and fair assessment. Machine Learning (ML), particularly through its branches such as deep learning and reinforcement learning, is revolutionizing CAT by enabling sophisticated analysis of large datasets, detailed behavior modeling, and flexible adaptation to diverse testing environments (Zhuang et al., 2022b). Despite the ML in CAT is still in its early stages, its potential is evident, with deep learning, natural language processing, and reinforcement learning showing significant promise in enhancing question characterization, scoring processes, and adaptive modeling (Mujtaba and Mahapatra, 2020). This survey aims to provide an overview and understanding of traditional statistical-based and recent ML-based CAT.
Application of CAT
CAT has transcended the boundaries of educational assessment, finding its place in diverse fields such as healthcare, sociology, and sports (Yurtcu and GÜZELLER, 2021).
  • Education. CAT, with its roots in the U.S. Department of Defense’s ASVAB test in the 1980s, has transformed education assessments. It personalizes tests for each individual, benefiting fields such as nursing licensing exams, graduate school exams (GRE and GMAT), and mathematics courses (Verschoor and Straetmans, 2010). There’s also a newer CAT variant called Multi-Stage Testing (MST). While traditional CAT adapts question by question, MST adapts in stages, with a set of questions being provided at each stage (Luecht and Nungester, 1998; Wainer and Kiely, 1987). MST is gaining popularity in exams like the GRE, law school exams (LSAC), and accounting exams (CPA).
  • Healthcare. In healthcare, particularly in mental health, there’s a crucial need for quick and accurate assessments. CAT meets this demand by offering quick evaluations akin to those conducted by skilled clinicians, thereby enhancing the overall efficiency of medical care. CAT for mental health assessments creates scales, including those for depression, anxiety, and suicide risk (Gibbons et al., 2018; Gibbons and deGruy, 2019; Gibbons et al., 2017, 2014), have shown promising results compared to traditional methods (Bagby et al., 2004).
  • Sociology. In the domain of sociology, CAT can be used in public opinion surveys. Montgomery and Cutler (2013) introduce an algorithm that adaptively selects survey questions. It not only enhances the accuracy of the survey results but also significantly reduces the number of questions needed, thereby streamlining the process and cutting down on associated costs.
  • Sports. The evaluation of specific skills is crucial for the development of athletes. Notably, Ando et al. (2018) innovatively apply CAT to the realm of football. They create a standardized assessment called the TSS-CAT, which utilizes video recordings of football tactics as test questions to gauge a player’s tactical proficiency. This groundbreaking assessment tool offers a tailored and efficient way for players, coaches, and managers to monitor and evaluate the progression of individual tactical skills throughout the training process.
3. Overview
An important assumption (Chang, 2015) of CAT is that examinee’s true proficiency level $\theta_{0}\in\mathbb{R}^{d}$ is constant throughout the test, where $d$ refers to the proficiency’s dimension; for example, $\theta$ could denote an overall ability level ($d=1$) or represent the degree of mastery across $d$ knowledge concepts. In CAT, the aim is to estimate examinees’ true proficiency levels (i.e., $\theta_{0}$) by having them answer questions. The ultimate objectives of CAT are twofold: (1) to use the responses to estimate an examinee’s proficiency $\theta$ such that it closely approximates the true proficiency $\theta_{0}$ by the end of the test, and (2) to select the most valuable and fitting questions for each examinee, thereby reducing test length.
Figure 2. The workflow of CAT: At step $t$, the selection algorithm adaptively selects next question $q_{t+1}$ based on examinee’s current proficiency ${\theta}^{t}$ estimated by CDM.
3.1. Task Formalization
As illustrated in Figure 2, at test step $t\in[1,2,...,T]$ in CAT, examinee’s current proficiency estimate $\hat{\theta}^{t}$ is estimated using previous $t$ responses; then leverage $\hat{\theta}^{t}$ to retrieve the next question $q_{t+1}$ from question bank $\mathcal{Q}$ to ask examinee, and receive his/her next response $y_{t+1}$. These interactions form a response sequence $\{(q_{1},y_{1}),(q_{2},y_{2}),...,(q_{T},y_{T})\}$, where $y_{t}=1$ if the response to $q_{t}$ is correct and 0 otherwise. To achieve the goals of CAT, the following two steps are necessary at each test step: Step 1: Proficiency Estimation. The Cognitive Diagnosis Model (CDM), denoted by $f(\cdot)$, acts as a user model, predicting the probability of a correct response by an examinee with proficiency $\theta$, which is denoted as $f(q,\theta)=P(y=1|q,\theta)$. The implementation of CDM often draws upon cognitive science (De La Torre, 2009) or psychometrics(Ackerman et al., 2003). To accurately estimate his/her proficiency at each step, various estimation methods can be used, e.g., Maximum Likelihood Estimation (MLE) or Bayesian Estimation. In applications, the binary cross-entropy loss is frequently utilized: at step $t$, given previous $t$ responses $X_{1:t}=\{(q_{1},y_{1}),(q_{2},y_{2}),...,(q_{t},y_{t})\}$, the corresponding empirical loss is:
(1) ${L}(\theta)=\sum_{(q,y)\in X_{1:t}}l(y,f(q,\theta))=-\sum_{(q,y)\in X_{1:t}}{y\log f(q,\theta)+(1-y)\log(1-f(q,\theta))},$
thus the current estimate of proficiency, $\hat{\theta}^{t}$, is obtained by minimizing the loss function $L(\theta)$: $\hat{\theta}^{t}=\mathop{\arg\min}_{\theta}{{L}(\theta)}$. Step 2: Question Selection. The heart of CAT is an algorithm that picks the next question $q_{t+1}$ from the question bank $\mathcal{Q}$, using examinee’s current proficiency estimate $\hat{\theta}^{t}$ as a guide:
(2) $q_{t+1}=\mathop{\arg\max}_{q\in\mathcal{Q}}\mathcal{V}_{q}(\hat{\theta}^{t}),$
where $\mathcal{V}_{q}(\hat{\theta}^{t})$ is the value of question $q$. For instance, $\mathcal{V}$ might be a measure of how much information the question will provide about the examinee’s proficiency, or it could be the output of a policy $\pi$ specifically designed to determine question selection. After receiving new response $y_{t+1}$, CDM updates and estimates proficiency $\hat{\theta}^{t+1}$. The above process will be repeated for $T$ times, ensuring the final step estimate $\hat{\theta}^{T}$ close to the true $\theta_{0}$, i.e.,
Definition 0 (Definition of CAT).
The goal of CAT is to find a question set $S=\{q_{1},q_{2},...,q_{T}\}$ of size $T$, such that the final step estimate $\hat{\theta}^{T}$, derived from $S$ and their corresponding response labels $y$, closely approximates the examinee’s true proficiency $\theta_{0}$:
(3) $\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|.$
However, solving this optimization problem directly is impractical, as the true proficiency $\theta_{0}$ is not observable, and even the examinees may not know their exact proficiency level. Consequently, existing methods are all approximations of this target. For instance, traditional statistical selection methods (Lord, 2012; Chang and Ying, 1996) utilize the asymptotic statistical properties of the MLE to reduce estimation uncertainty, such as choosing questions that closely match the examinee’s current estimated proficiency $\hat{\theta}^{t}$. More recent Subset Selection approaches (Zhuang et al., 2023) try to identify a theoretical approximation of $\theta_{0}$ to serve as a new objective for optimization. For further details, refer to Section 5. Meanwhile, as a practical system, considerations extend beyond this proficiency estimation objective (Definition 1). Factors such as CAT’s question exposure control, robustness, fairness, and search efficiency must be addressed as well. An exhaustive discussion of these factors is presented in Section 7. Evaluation Methods: To validate the accuracy of the estimated proficiency, two primary approaches are employed: 1) using the examinee’s estimated values within the CDM to predict the correctness label $y$ of the responses on his/her reserved response data, often measured by cross-entropy; 2) using simulation to generate true proficiency values $\theta_{0}$, simulating the examinee’s responses to each question. Then the Mean Squared Error (MSE) between the estimates and the simulated true values can be calculated. The details can be found in Section 8.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=15em,font=,, where level=3text width=17em,font=,, [ Computerized Adaptive Testing, ver [ Cognitive Diagnosis Model
     Section 4, fill=red!10, text width=15em [ Latent Trait Model (4.1), fill=red!10, text width=19em [  IRT (Embretson and Reise, 2013; Hambleton et al., 1991; Lord, 2012), MIRT (Reckase, 2006), VI-IRT (Wu et al., 2020)., leaf, text width=22em ] ] [ Diagnostic Classification Model (4.2), fill=red!10, text width=19em [  DINA (De La Torre, 2009; Von Davier, 2014), G-DINA (de la Torre, 2011),
 AHM (Leighton et al., 2004), FuzzyCDF (Liu et al., 2018)., leaf, text width=22em ] ] [ Deep Learning Model (4.3), fill=red!10, text width=19em [  DIRT (Cheng et al., 2019), NeuralCD (Wang et al., 2020c), IRR (Tong et al., 2021),
 RCD (Gao et al., 2021), HierCDF (Li et al., 2022), DeepCDF (Gao et al., 2022),
 BETA-CD (Bi et al., 2023), AGCDM (Pei et al., 2022), SCD (Wang et al., 2023b).,leaf, text width=22em ] ] ] [ Selection Algorithm
   Section 5, fill=blue!10, text width=15em [ Statistical Algorithms (5.1), fill=blue!10, text width=18em [  Classic Information: FIsher Info (Lord, 2012), KL Info (Chang and Ying, 1996; Chang, 2015);
 Additional Information: MPWI (Barrada et al., 2009a), MEI (van der Linden, 1998), thOpt (Barrada et al., 2006);
 CD-CAT: KL algorithm (Cheng, 2009), Shannon entropy (Tatsuoka, 2002);
 Dual-objective CD-CAT (Kang et al., 2017; Zheng et al., 2018).,leaf, text width=27em ] ] [ Active Learning Algorithms (5.2), fill=blue!10, text width=18em [  Uncertainty method (Veldkamp and Verschoor, 2019), Expected Model Change (Bi et al., 2020)., leaf, text width=27em ] ] [ Reinforcement Learning (5.3), fill=blue!10, text width=18em [  SSP (Gilavert and Freire, 2022), POMDP (Nurakhmetov, 2019; Gilavert and Freire, 2022; Chen et al., 2018), DQN (Li et al., 2023),
 GMOCAT (Wang et al., 2023a), NCAT (Zhuang et al., 2022b)., leaf, text width=27em ] ] [ Meta Learning Algorithms (5.4), fill=blue!10, text width=18em [  BOBCAT (Ghosh and Lan, 2021; Feng et al., 2023), DL-CAT (Ma et al., 2023), SACAT (Yu et al., 2023)., leaf, text width=27em ] ] [ Subset Selection Algorithms (5.5), fill=blue!10, text width=18em [  BECAT (Zhuang et al., 2023), SEM (Mujtaba and Mahapatra, 2021)., leaf, text width=27em ] ] ] [ Question Bank Construction
     Section 6, fill=yellow!10, text width=15em [ Question Characteristics Analysis (6.1), fill=yellow!10, text width=20em [  Expert-based (López-Cuadrado et al., 2008; Kozierkiewicz-Hetmańska and Poniatowski, 2014);
 Statistic-based (Chen et al., 2004; Magno, 2009; Chang and Yang, 2009; Sun et al., 2014; Xiong et al., 2022);
 Deep Learning: CNN (Huang et al., 2017), RNN (Qiu et al., 2019; Huang et al., 2019; Lei et al., 2021),
 NLP (Yin et al., 2019; Ning et al., 2023). , leaf, text width=22em ] ] [ Question Bank Development (6.2), fill=yellow!10, text width=20em [  Expert-based (Reckase, 2010; He and Reckase, 2014; Van Der Linden et al., 2000; Way et al., 2005; van der Linden et al., 2006; Ariel et al., 2004; Gulliksen, 2013; Swanson and Stocking, 1993)., leaf, text width=22em ] ] ] [ Test Control
  Section 7, fill=green!10, text width=15em [ Content Control (7.1), fill=green!10, text width=13em [  Sympson-Hetter (Sympson and Hetter, 1985; van der Linden and Veldkamp, 2004, 2007; Barrada et al., 2009b);
 A-Stratified (Chang and Ying, 1999; Chang et al., 2001; Yi and Chang, 2001; Barrada et al., 2006, 2014)., leaf, text width=20em ] ] [ Fairness (7.2), fill=green!10, text width=13em [  Bias in CDMs (Thompson, [n.d.]; Kizilcec and Lee, 2022);
 Bias in Question Bank (Hambleton et al., 1991; Chu and Lai, 2013; Camilli and Shepard, 1994; Mellenbergh, 1989; Roberts, [n.d.]);
 Bias in Selection Algorithms (Lord, 2012);
 Equating (Brigman and Bashaw, 1976; van der Linden, 2000; JAN, [n.d.]; Sawaki, 2001).,leaf, text width=20em ] ] [ Robustness (7.3), fill=green!10, text width=13em [  Robust CAT(Veldkamp and Verschoor, 2019; Zhuang et al., 2022a).,leaf, text width=20em ] ] [ Search Efficiency (7.4), fill=green!10, text width=13em [  PSO (Huang et al., 2009), SECAT (Hong et al., 2023).,leaf, text width=20em ] ] ] ]
Figure 3. Summary of Computerized Adaptive Testing methods in machine learning perspective.
3.2. Categorization
As shown in Figure  3, we categorize existing CAT work into four parts based on the above three main components and the test control in CAT applications: 1) Cognitive Diagnostic Model, 2) Selection Algorithm, 3) Question Bank Construction, and 4) Test Control. Each part is further divided based on the different techniques employed. The following four sections (Section 4-7) will provide separate introductions and reviews of the relevant research within these four categories.
4. Cognitive Diagnosis Model
The existing methods for CDM can be categorized into three main types: Latent Trait Model, Diagnostic Classification Model, and Deep Learning Model. Such classification is grounded on the nature of proficiency representation ($\theta$) in CAT—ranging from an overall numerical ability value (Latent Trait Model), to discrete cognitive states across different knowledge concepts (Diagnostic Classification Model), to a unified modeling approach via deep learning techniques (Deep Learning Model). The choice of model should depend on the specific goals of the assessment, the nature of the data, and the resources available. Regardless of the chosen CDM for estimating proficiency, the objective remains consistent: to minimize the error between the estimate and the true value at each step, expressed as $\|{\hat{\theta}^{t}}-\theta_{0}\|\to 0$.
4.1. Latent Trait Model
In latent trait models, the examinee’s proficiency is handled as overall ability. Item Response Theory (IRT) (Embretson and Reise, 2013) is a representative work. In IRT, the examinee’s proficiency is modeled as scalars, which can be denoted as $\theta$. One of the most widely used models in IRT is the Three-Parameter Logistic Model (3PL-IRT). It utilizes a logistic-like interaction function to model the probability of examinee’s correct response to question $j$, i.e.,
(4) $f(q_{j},\theta)=c_{j}+\frac{1-c_{j}}{1+e^{-\alpha_{j}(\theta-\beta_{j})}}.$
The 3PL-IRT model introduces three parameters ($\beta_{j},\alpha_{j},c_{j}$) for each test question $j$: The difficulty parameter $\beta_{j}$ corresponds to the level of proficiency at which an examinee has a 50% chance of answering the question correctly; The discrimination parameter $\alpha_{j}$ describes how well the question differentiates between examinees with different ability; The guessing parameter $c_{j}$ represents the probability that an examinee with a very low proficiency will answer the question correctly. In CAT systems, these parameters are pre-calibrated and remain fixed during the testing process, with annotation and calibration methods detailed in Section 6. IRT takes into account the number of questions answered correctly and the difficulty of the question. Today, almost all major adaptive tests, such as SAT and GRE, are developed by using IRT, because the methodology can significantly improve measurement reliability and interpretability (An and Yung, 2014). Multidimensional IRT (MIRT) (Reckase, 2006), on the other hand, extends IRT to multiple dimensions, allowing for the modeling of multiple latent traits simultaneously. Despite the great interpretability of (M)IRT models, their performance is constrained by the simplicity of the interaction function, and they lack fine-grained modeling about examinee’s cognitive states on individual knowledge concepts.
4.2. Diagnostic Classification Model
Diagnostic Classification Model (DCM) is another representative class of CDMs, focusing on discrete knowledge concepts. Specifically, in DCMs, examinee proficiency is knowledge concept-wise and usually dichotomous, which indicates whether an examinee has mastered a knowledge concept or not. The DINA method (De La Torre, 2009; Von Davier, 2014) models examinee proficiency $\theta=\{\theta_{(1)},\theta_{(2)},...,\theta_{(K)}\}$ as their dichotomous knowledge mastery levels on all $K$ concepts. Given the Q-matrix $Q\in\mathbb{R}^{|\mathcal{Q}|\times K}$which is a binary matrix that indicates which knowledge concepts are associated with a question in bank $\mathcal{Q}$. DINA method focuses only on the knowledge concepts related to the target question $j$, where $Q_{jk}=1$. Thus, the examinee’s binary response variable (with proficiency $\theta$) to question $j$ is $\prod_{k,Q_{jk}=1}\theta_{(k)}$, and models questions as “slip” and “guess” parameters:
(5) $f(q_{j},\theta)=(1-s_{j})^{\prod_{k,Q_{jk}=1}\theta_{(k)}}g_{j}^{1-\prod_{k,Q_{jk}=1}\theta_{(k)}},$
where $s_{j}$ is the slip parameter, indicating the likelihood of an incorrect response despite mastery, and $g_{j}$ is the guess parameter, reflecting the chance of a correct guess in the absence of mastery. Its extension G-DINA(de la Torre, 2011) provides a granular view of examinee proficiency, while FuzzyCDF(Liu et al., 2018) leverages fuzzy set theory for nuanced diagnostics from both objective and subjective data. Another approach, the Attribute Hierarchy Method (AHM)(Leighton et al., 2004), applies rule space theory to structure knowledge dependencies and align examinee proficiencies with the nearest ideal cognitive patterns to obtain diagnostic results. Compared to Latent Trait Models, Diagnostic Classification Models offer a more granular and comprehensive assessment of examinee proficiencies. They are particularly adept at providing detailed feedback on an individual’s strengths and weaknesses across multiple knowledge concepts, which is important for CAT and targeted further interventions. These models underscore a critical shift towards a more nuanced understanding of learning and proficiency, recognizing the multifaceted nature of knowledge acquisition and adaptive testing.
4.3. Deep Learning Model
In recent years, the rapid growth of deep learning techniques stimulates the development of deep learning-driven CDMs. Compared to traditional models, deep learning methods are more suitable for cognitive diagnosis in large-scale data scenarios (e.g., online learning platforms) due to their efficiency and ability to learn the complex interaction pattern between examinees and questions. For instance, DIRT (Cheng et al., 2019) uses a neural network to capture semantic information from question texts to empower accuracy. NeuralCD (Wang et al., 2020c) utilizes a non-negative full connection neural network to capture the complex interaction, with the ability to generalize to other CDM methods. Considering the complex heterogeneous relationships between examinees, questions, and knowledge concepts, massive efforts have also been made to leverage them to enhance cognitive diagnosis (Gao et al., 2021; Li et al., 2022; Gao et al., 2022).
Discussion 1.
In the CAT process, only a limited number of examinee responses can be obtained for proficiency estimation, due to the adaptive nature of the test. To some extent, CAT can be viewed as a cognitive diagnosis under a cold start scenario. The performance of the CDM is a critical factor in ensuring the accuracy of proficiency estimations within CAT. Meanwhile, it is important to note that the choice of the CDM can significantly influence the selection of corresponding question selection algorithm.
5. Selection Algorithm
The selection algorithm is CAT’s core of implementing adaptivity and is the focal point of this survey. It utilizes the proficiency estimate obtained from the CDM (introduced in the above section) to choose the next most suitable question, ensuring an accurate estimation of proficiency while using the fewest possible questions. Question selection algorithms can be categorized into traditional methods based on statistical information, as well as more recent machine learning methods, e.g., data-driven approaches (i.e., Reinforcement Learning and Meta Learning), and Subset Selection are becoming increasingly prevalent.
5.1. Statistical Algorithms
In this subsection, we first establish the general framework and methodology for statistical question selection, which typically involves defining a specific information metric, e.g., Fisher information, or Kullback-Leibler information.
5.1.1. General Framework
In general, a practical approach to designing a selection algorithm involves developing quantitative methods to assign a numerical value to each question in the bank $\mathcal{Q}$. The computer can then use this value assignment as a heuristic function to continuously select the most valuable questions during the personalized testing process, achieving adaptive question selection. In a more systematic view, well-adopted by classical question selection strategies, is from a statistical perspective, defining the value of a question as the informativeness it provides about the examinee’s potential ability estimation. The next question index $j_{t+1}$ can be selected from bank $\mathcal{Q}$ based on current estimate $\hat{\theta}^{t}$:
(6) $j_{t+1}=\arg\max_{q_{j}\in\mathcal{Q}}\mathcal{I}_{j}(\hat{\theta}^{t}),$
where $\mathcal{I}_{j}(\cdot)$ is the informativeness of question $q_{j}$ (e.g., Fisher information). As illustrate in Definition 1, CAT assumes that each examinee has a true proficiency value ($\theta_{0}$) and it is considered as a parameter estimation process. The informativeness of a question can thus be interpreted as the expected contribution of the response on this question to the parameter estimation. This concept will be reflected in various selection algorithms discussed later. Fisher Information. In the parameter estimation problems, Fisher Information (Rissanen, 1996) is a concept from information theory and statistics that measures the amount of information that an observable random variable carries about an unknown parameter upon which the probability of that random variable depends. In the context of CAT, Fisher Information is often used to quantify the amount of information that a question provides about an examinee’s proficiency (Cheng, 2008). Specifically, we consider a random variable $X_{j}=(q_{j},y_{j})$ for which the pdf or pmf is $f(q_{j},\theta)$, where $\theta$ is the unknown parameter. The fisher info contained in the random variable $X_{j}$ is defined as: $\mathcal{I}_{j}(\theta)=\mathbb{E}_{y_{j}}[(\nabla_{\theta}L(X_{j}|\theta))^{2}]=\frac{(\nabla_{\theta}f(q_{j},\theta))^{2}}{f(q_{j},\theta)(1-f(q_{j},\theta))}$, where $L(X|\theta)={y\log f(q,\theta)+(1-y)\log(1-f(q,\theta))}$ is the likelihood function of $X$ with respect to the parameter $\theta$. Thus, when using 3PL-IRT to model the $f(q,\theta)$ and given current estimate $\hat{\theta}^{t}$, the Fisher Information of question $j$ can be calculated as:
$\mathcal{I}_{j}(\hat{\theta}^{t})=\frac{(1-c_{j})\alpha_{j}^{2}e^{-\alpha_{j}(\hat{\theta}^{t}-\beta_{j})}}{(1+e^{-\alpha_{j}(\hat{\theta}^{t}-\beta_{j})})^{2}[1-c_{j}+c_{j}(1+e^{-\alpha_{j}(\hat{\theta}^{t}-\beta_{j})})]}.$
One crucial property of Fisher information is that its reciprocal (matrix inverse), is the variance (covariance matrix) of the asymptotic distribution of the maximum likelihood estimate:
Theorem 1 (The asymptotic distribution of MLE proficiency estimate (Ross, 2014)).
At each step $t$, based on the observation of examinee’s previous $t$ responses, the current proficiency estimate $\hat{\theta}^{t}$ (estimated by MLE) satisfies the asymptotic normal distribution: $\hat{\theta}^{t}\sim\mathcal{N}\left(\theta_{0},\frac{1}{t\mathcal{I}(\theta_{0})}\right).$
Obviously, it can be obtained that as the number of test questions ($t$) or the Fisher information ($\mathcal{I}$) increases, the variance decreases. Since the estimated value is asymptotically unbiased (i.e., its mean equals the true value $\theta_{0}$), a decrease in its variance results in the distribution becoming more “tight” (higher $p(\hat{\theta}^{t}=\theta_{0})$), thereby reducing the uncertainty of the estimated proficiency $\hat{\theta}^{t}$. Due to the unknown $\theta_{0}$, maximizing the Fisher Information (MFI) at the current estimate $\hat{\theta}^{t}$ essentially serves to lower the variance of the proficiency estimate and reduce uncertainty, thereby improving the estimation efficiency. MFI has been widely popular in the development of personalized testing over the decades and extensively applied in practice as well. However, the effectiveness of MFI is most pronounced when the estimated value is close to the unknown true value $\theta_{0}$ (Chang, 2015). Therefore, using the current estimate in place of $\theta_{0}$ can sometimes lead to inaccuracies if the current estimate is far from the true value. Fisher Information is a mathematical concept that exists independently of CAT. Notably, this concept finds a natural extension in Multidimensional IRT (MIRT), where it transitions from a real number to a matrix (Hooker et al., 2009), facilitating the assessment of multiple abilities or traits simultaneously. Specifically, the information matrix provided by question $j$ at proficiency $\theta$ (now a vector) is defined as: $\mathcal{I}_{j}(\theta)=\mathbb{E}_{y_{j}}\left[\nabla\log L(X_{j}|\theta)\nabla\log L(X_{j}|\theta)^{T}\right]$. This Fisher Information matrix is instrumental in MIRT-based personalized assessment systems, as its inverse corresponds to the covariance matrix for maximum likelihood proficiency estimates. Researchers have exploited this relationship to develop many question selection strategies such as D-Optimization, A-Optimization, and E-Optimization, which respectively focus on maximizing the matrix’s determinant, minimizing the trace of its inverse, and maximizing its minimum eigenvalue (Reckase, 2006). These strategies, complemented by Bayesian optimization variants, have considerably advanced the statistical techniques available for question selection in adaptive testing. Kullback-Leibler Information. Fisher information is a commonly used measure in CAT due to its theoretical foundation, simplicity in mathematical form, and computational efficiency. However, the application of Fisher information might not be as efficient as expected if the estimators are not close to the true value $\theta_{0}$ (Chang, 2015). This issue becomes particularly evident during test’s initial stage, when the test length (number of questions) is too short to provide an accurate proficiency estimate. Consequently, relying on the MSI may not yield a satisfactory results, because Fisher information, often referred to as local information, only considers the information at a specific point ($\hat{\theta}^{t}$). To address this issue, Chang and Ying (1996) proposed a global information measure based on Kullback-Leibler (KL) divergence, also known as KL information: Let $\theta_{0}$ be the true proficiency, for any $\theta$, the KL information of the question $q_{j}$ (with response $X_{j}$) which measures the discrepancy between the two probability distributions specified by $\theta_{0}$ and $\theta$, is defined by
$KL_{j}(\theta||{\theta}_{0})=\mathbb{E}_{y_{j}}\log\left[\frac{L(X_{j}|\theta_{0})}{L(X_{j}|\theta)}\right]=f(q_{j},\theta_{0})\log\left[\frac{f(q_{j},\theta_{0})}{f(q_{j},\theta)}\right]+(1-f(q_{j},\theta_{0}))\log\left[\frac{1-f(q_{j},\theta_{0})}{1-f(q_{j},\theta)}\right].$
The important feature is that KL is a function of two levels, $\theta$ and $\theta_{0}$. $KL(\theta||{\theta}_{0})$ represents the discrimination power of the question on the two levels and does not require that $\theta$ be close to $\theta_{0}$. The corresponding algorithm for question selection that maximizes KL information involves integrating over $\theta$ in Eq.(5.1.1) and substituting the current estimate for $\theta_{0}$:
(7) $\mathcal{I}_{j}(\hat{\theta}^{t})=\int_{\hat{\theta}^{t}-\delta}^{\hat{\theta}^{t}+\delta}KL_{j}(\theta||\hat{\theta}^{t})d\theta.$
where $\delta=3/\sqrt{t}$ and the integration range is relatively large at the test’s outset and gradually narrows as $t$ increases. In MIRT, the global KL divergence information extends from a univariate integral to a multivariate integral.
Figure 4. Illustration of KL and Fisher information functions for two questions (Question 1: $\alpha=1.7,\beta=1.9,c=0.1$; Question 2: $\alpha=2.3,\beta=0.5,c=0.3$). Assuming the current proficiency estimate $\hat{\theta}^{t}=1$. The KL information (left) for the given question represents an integral centered around $\hat{\theta}^{t}$, while the Fisher information (right) corresponds to the value at the specific point $\hat{\theta}^{t}$.
Essentially, the KL method identifies questions that will provide the greatest differentiation between the examinee’s possible proficiency levels. Figure 4 illustrates the KL and Fisher information functions for two distinct questions. The information conveyed by KL Information is not confined to a single point value like Fisher; instead, it spans a wide range of proficiency levels $\theta$. This is why KL information is global while Fisher information is local (Chang, 2015). Interestingly, for $\theta$ near $\theta_{0}$, KL Information and Fisher information are always close. If we envision KL Information as a curve, Fisher information corresponds to its curvature (second derivative) at $\theta=\theta_{0}$. This suggests that Fisher information can be derived from KL Information, but the converse is not true.
5.1.2. Advances and Generalization
[…]
Discussion 2.
Selection algorithms in CAT have primarily relied on the above statistical heuristic approaches, which require domain experts to consider every possible testing scenario and manually design corresponding selection algorithms. These methods are model-specific, requiring distinct selection strategies for different CDMs. For example, the above MFI strategy (Lord, 2012) is specifically crafted for (M)IRT. Consequently, previous statistical methods lack flexibility and are labor-intensive, and the selection algorithm must be re-designed if the underlying CDM changes.
5.2. Active Learning Algorithms
An ideal CAT framework should be model-agnostic, meaning that the selection algorithms can adapt to different CDMs while remaining effective. Researchers are exploring a broader approach to the CAT problem by investigating its correlation with a versatile machine learning technique—Active Learning (Krishnakumar, 2007). Active Learning is to choose valuable data actively, thus can train better models with less data. This method has improved data efficiency and cut annotation costs in numerous learning tasks (Budd et al., 2021). As shown in Figure 5, in active learning, each cycle begins with a learning model and a strategy for selecting data. The system selects an unlabeled data (set) to label at each iteration, using this to supplement the limited labeled data and improve the model’s performance. The key point behind this is how to design the data selection strategy. Existing methods typically view selection from two perspectives: informativeness and representativeness. Informativeness-based algorithms choose samples that significantly reduce a learning model’s uncertainty (Yoo and Kweon, 2019). Representativeness-based algorithms, on the other hand, choose samples that best reflect the overall patterns of unlabeled data (Ghorbani et al., 2022; Wang et al., 2020b), and their combinations (Wang et al., 2020b).
Figure 5. Top: The Active Learning Framework. Bottom: The relationship and correspondence between each component of Active Learning and those of CAT.
The idea and the pattern of active learning are similar to that of CAT, inspiring a general framework. To be specific, in CAT systems, the CDM serves as the learning model, the question selection algorithm serves as the data selection strategy, and the examinee’s question-taking behavior can be regarded as data annotation. The goal is to select as few samples (questions) as possible to effectively train the learning model (estimate the proficiency parameters in CDM). This framework allows for a model-agnostic perspective in designing CAT solutions, as it doesn’t require the consideration of the CDM’s specific implementation details. For example, Bi et al. (2020) propose a model-agnostic adaptive testing framework along these lines. Its core idea is that the change (i.e., the gradient in machine learning) in the proficiency estimate parameter, upon observing an examinee’s response to a question,
(8) $j_{t+1}=\arg\max_{q_{j}\in\mathcal{Q}}\mathbb{E}_{y_{j}}\left\|\nabla_{\theta}L(X_{1:t}\cup\{(q_{j},y_{j})\}|\theta)\right\|.$
It calculates the expected norm of each resulting gradient vector, taking into account the unknown response label $y_{j}$ during the selection process. The intuition behind this framework is that it prefers questions that are likely to most influence the proficiency estimation (i.e., have the greatest impact on its parameters), regardless of the resulting label. The active learning technique successfully alleviates the model-specific issues in CAT.
Discussion 3.
The above heuristic rules-based approaches (i.e., statistical algorithms and Active Learning algorithms) cannot be improved by learning from large-scale examinee response data (Ghosh and Lan, 2021; Li, 2020). To overcome this limitations and enable scalable personalized testing, data-driven approaches have gained prominence. These approaches using machine learning and deep learning can achieve better performance. They leverage techniques such as Reinforcement Learning (Section 5.3) and Meta Learning (Section 5.4) to automatically make suitable question selection by learning patterns from large-scale examinee response data.
5.3. Reinforcement Learning Methods
Reinforcement learning (RL), a subfield of machine learning, is a powerful approach that enables an agent to learn how to make optimal decisions automatically (Sutton and Barto, 2018). It has been successfully applied in various domains, including robotics, autonomous vehicles, education, and healthcare (Arulkumaran et al., 2017; Kober et al., 2013). In RL, an agent interacts with an environment and receives feedback in the form of rewards or penalties based on its actions. The goal is to learn a policy $\pi$, which can maximize the long-term cumulative reward. As shown in Figure 6, the policy can be learned by exploring the environment and learning from its consequences of actions. In essence, researchers in CAT utilize RL methodologies to address a question: Can the selection algorithm (policy) be automatically learned and optimized from data or student interactions, thus circumventing the necessity for expert intervention? Markov Decision Process Formulation. The interaction between agent and environment can be viewed as a Markov Decision Process (MDP) (Feinberg and Shwartz, 2012). Specifically, at each step, the agent observes current environment’s state ($s$), and interacts with the environment by selecting its actions ($a$). Simultaneously, the agent receives a reward ($r$) from these interactions, influencing or changing the current state of the environment. The objective is to select a best sequence of actions, resulting in the highest cumulative reward ($\sum r$). Therefore, most RL problems are formally described as estimating the optimality of the agent’s behavior in a given state (value-based methods (Sutton and Barto, 2018)) or the optimality of the action policy itself (policy-based methods (Agarwal et al., 2021)) or the hybrid approaches (Mnih et al., 2016). We next formulate the CAT problem as an MDP. Here are the definitions of the key elements in RL within an adaptive learning/testing system and the overall RL framework of CAT is in Figure 6:
Figure 6. The overall Reinforcement Learning framework of CAT. The objective is to optimize the selection algorithm $\pi$ (i.e., policy) by exploring the large-scale examinee response data (i.e., environment).
  • State: A state $s_{t}\in S$ represents the current condition or situation at each test step $t$. It captures relevant information about examinee and the CAT system. Generally, the state includes the examinee’s previous response sequence (or a latent vector to represent the current proficiency estimate (Li et al., 2020)) and the candidate questions in the question bank (Zhuang et al., 2022b; Nurakhmetov, 2019): $s_{t}=(\{q_{1},y_{1},...,q_{t},y_{t}\},\mathcal{Q})$.
  • Action: An action $a$ refers to the choices that the CAT system can take in current state $s_{t}$, i.e., the selection of the next question from bank $q_{t+1}\in\mathcal{Q}$.
  • Transition: The transition function is the probability of seeing state $s_{t+1}$ after taking action $q_{t}$ at current state $s_{t}$: $P(s_{t+1}|s_{t},q_{t+1})$. At each step, the uncertainty comes from the examinee’s response correctness label $y_{t+1}$ to question $q_{t+1}$.
  • Reward: A reward $r$ is a scalar feedback that the CAT receives after selecting a question for the examinee at each step. To achieve CAT’s goal (i.e., $\min{\|\hat{\theta}^{T}-\theta_{0}\|}$ in Definition 1), the reward function can be defined as the accuracy of proficiency estimation at each step [^1] As the true value is often unobtainable, it is commonly derived through simulation experiments. (Ma et al., 2023; Nurakhmetov, 2019; Wang et al., 2023a), i.e., $\|\hat{\theta}^{t}-\theta_{0}\|$, or the performance prediction loss of $\hat{\theta}^{t}$ on the held-out response data (Zhuang et al., 2022b; Shin and Bulut, 2022). This reward signal is pivotal in guiding the policy $\pi$ to select the best-fitting question that can reduce the estimation error.
Recently, with the advancements in deep learning, an increasing number of studies are leveraging Deep Reinforcement Learning (DRL) to tackle the MDP problems in CAT. Li et al. (2023) utilize the Deep Q-Network (DQN) to represent the action-value function $Q_{w}(s,q)$, representing the value of choosing question $q$ in state $s$, and $w$ denotes its parameter of the network’s fully connected layer. The most suitable question is selected according to the policy:
(9) $\pi^{*}(q|s)=\arg\max_{q\in\mathcal{Q}}{Q}_{w}(s,q).$
For modeling the complex interactions between examinees and questions during the practical test, Zhuang et al. (2022b) recently design a transformer-based Q-Network, which can identify and extract the perturbations in examinee performance (e.g., guess and slip factors). Stochastic Shortest Path Formulation. Furthermore, CAT can be defined as a Stochastic Shortest Path (SSP) problem (Bertsekas and Tsitsiklis, 1991), which is a special case of MDP. In an SSP, the objective is to find the shortest path (i.e., the minimum test step) from a given initial state $s_{0}$ to goal states. In CAT, the goal state typically represents the completion of the test or the attainment of a predetermined level of proficiency estimation precision. The approach by Gilavert and Freire (2022) uses Linear Programming to find the optimal testing policy $\pi^{*}$, treating the problem like a flow network where each state must have balanced inflow and outflow, except for the start and end points. It denotes variables $x_{s,a}$ as the expected accumulated occurrence frequency for every pair (state $s\in S$, action $q\in\mathcal{Q}$), and equalizes $in(s)$ and $out(s)$ flow model for every state $s$. The flow into a state $s$ is the sum of the expected frequencies of all actions in all other states $s^{\prime}$ that lead to $s$: ${in}(s)=\sum_{s^{\prime},q}x_{s^{\prime},q}P(s|s^{\prime},q)$. The flow out of a state $s$ is the sum of the expected frequencies of all actions in state $s$: ${out}(s)=\sum_{q}x_{s,q}$. The objective function is to maximize the total expected reward $r$, which is the sum of the expected frequencies times the immediate rewards $r(s,q)$ for all state-action pairs: $\min_{x_{s,q}}\sum_{s\in S,q\in\mathcal{Q}}{x_{s,q}r(s,q)}$. Thus the optimal policy $\pi^{*}$ can be obtained by:
(10) $\pi^{*}(q|s)=\frac{x_{s,q}}{\sum_{q^{\prime}\in\mathcal{Q}}x_{s,q^{\prime}}}.$
Partial-Observable MDP Formulation. Partial-Observable MDP (POMDP) is a generalization of MDP to handle situations where the agent does not fully observe the environment (Hoerger and Kurniawati, 2021). Previously, it assumes that the proficiency of an examinee can be perfectly inferred from their previous responses, then CAT can be modeled as an MDP, where the proficiency estimate can be the states. However, in a more general scenario, the proficiency cannot be perfectly inferred from their previous responses. Because there is always some uncertainty involved (Zhuang et al., 2022a; Drousiotis et al., 2021):
  1. (1)
    Perturbations in Responses: The assessment process is subject to various perturbations stemming from examinee behavior and the testing environment. For example, an examinee might guess an answer correctly or misunderstand a question they would typically answer correctly. This variability introduces uncertainty and error in estimation.
  2. (2)
    Limited Information: Particularly at the onset of an assessment, the test has only sampled a small portion of the examinee’s knowledge through limited questions. Consequently, the responses may not be entirely indicative of the examinee’s comprehensive proficiency.
  3. (3)
    Complexity of “proficiency”: Proficiency isn’t a simple attribute. It can encompass various aspects like understanding, analysis, and application of all the knowledge concepts. Responses to a constrained set of questions may fall short of capturing the full breadth of these dimensions, potentially leading to an incomplete portrayal of an examinee’s proficiency.
To this end, many works (Nurakhmetov, 2019; Gilavert and Freire, 2022; Chen et al., 2018) model CAT as a POMDP. Compared with MDP, the POMDP model has two additional elements. $O$: A set of observations; $Z$: Observation probabilities. $Z(o|s^{\prime},q)$ is the probability of making observation $o$ after selecting question $q$ and transitioning to state $s^{\prime}$. The states can still represent the proficiency of the examinees, but now these states are not fully observable. Instead, these works maintain a probability distribution $b(s)$, called the belief state, over the possible proficiencies of the examinee given his/her previous responses. The belief state is typically updated using Bayes’ rule. When the agent select action (question) $q$ in belief state $b$ and makes observation $o$, it updates its belief state to $b^{\prime}(s^{\prime})$: $b^{\prime}(s^{\prime})=\eta{Z(o|s^{\prime},q)\sum_{s\in S}P(s^{\prime}|s,q)b(s)}$. where $\eta$ is a normalizing constant. POMDPs can be solved by many algorithms, such as Grid-based algorithms (Hoerger and Kurniawati, 2021), Monte Carlo tree search (Schwartz et al., 2022). However, due to its partial observability of the environment, POMDPs are more challenging to solve than MDPs.
5.4. Meta Learning Algorithms
Another data-driven machine learning approach that can address this complex CAT problem is meta-learning (Finn et al., 2017): It involves training a model on various tasks to acquire cross-task knowledge or learn how to learn efficiently. Specifically, the base-learner is trained across a variety of related tasks, allowing it to gather cross-task insights and general knowledge about how to learn efficiently. Then, the meta-learner leverages this knowledge to swiftly adapt to new, unseen tasks (Huisman et al., 2021). In CAT, each examinee’s testing process can be seen as a task because it involves selecting appropriate test questions based on his/her own proficiency level. The selection algorithm can be regarded as a form of general knowledge because it represents the accumulated knowledge and experience gained from a diverse set of examinees (Figure 7(right)). This knowledge can include the best policy for question selection, information about the characteristics of different test questions, the examinee proficiency prior, etc. By learning from these diverse examinees (tasks) in the large-scale response dataset, it can acquire a good question selection that can adapt to individual examinees.
Figure 7. The overall Meta Learning framework of CAT, and this figure is adapted from (Ghosh and Lan, 2021). The objective is to optimize the selection algorithm $\pi$ by exploring the large-scale examinee response data.
Bi-Level Optimization. One particular class of meta-learning methods that has shown promise is bi-level optimization, which involves optimizing both the global model’s parameters and the task-specific parameters simultaneously. This approach allows the model to learn how to adapt its learning process to different tasks, resulting in improved performance and generalization. Ghosh and Lan (2021) proposes a bi-level optimization framework for CAT (BOBCAT) to directly learn the data-driven selection algorithm $\pi$. Specifically: let $N$ denote the number of examinees in the response dataset for training $\pi$. The responses of each examinee $i$ are randomly divided into a support set $\mathcal{D}_{s}^{i}$ and a query set $\mathcal{D}_{u}^{i}$, where $\pi$ sequentially select a total of $t$ questions $\{q_{1},...,q_{t}\}$ from $\mathcal{D}_{s}^{i}$, observe their responses, and predict their response on the held-out query set $\mathcal{D}_{u}^{i}$. The global knowledge (i.e., selection algorithm $\pi$ and global parameters $\gamma$) is redefined as the objective of bi-level optimization:
(11) $\min_{\pi,\gamma}\frac{1}{N}\sum_{i=1}^{N}{\sum_{(q,y)\in\mathcal{D}_{u}^{i}}{l(y,f(q,\hat{\theta}_{i}))}},$
(12) $\mathrm{s.t.}\;\;\hat{\theta}_{i}=\mathop{\arg\min}_{\theta_{i}}{\sum_{(q,y)\in\mathcal{D}_{s}^{i}}{l\left(y,f\left(q,\theta_{i}\right)\right)}},$
(13) $\mathrm{where}\quad q_{t+1}\sim\pi\left(q|q_{1},y_{i(1)},...,q_{t},y_{i(t)}\right)\in\mathcal{D}_{s}^{i}.$
Figure 7(left) shows the overall meta learning framework. In the inner-level (Eq.(12)), the question in the support set $\mathcal{D}_{s}^{i}$ for examinee $i$ is sequentially selected by $\pi$, according to his/her previous responses; then binary cross-entropy loss $l(\cdot)$ on $\mathcal{D}_{s}^{i}$ is minimized for estimating the proficiency $\hat{\theta}_{i}$ for the outer-level. In the outer-level (Eq.(11)), the loss of the estimate $\hat{\theta}_{i}$ on the query set $\mathcal{D}_{u}^{i}$ is minimized to learn the selection algorithm $\pi$ and the global parameters $\gamma$. The global parameters $\gamma$ refer to the question characteristics and the examinee ability prior. The algorithm $\pi$ is also model-agnostic. It could be adapted to the given CDM ($f$) automatically by optimizing this problem for efficient selection. Once the question selection algorithm is trained, its parameters do not update during the CAT process and adaptively selects the next item based on previous response behaviors. Based on the above bi-level definition, there have been increasing efforts to improve upon it. Ma et al. (2023) propose a flexible optimization framework Decoupled Learning CAT (DL-CAT). The original BOBCAT obtains the parameters of two modules (i.e., examinee proficiency estimation and question selection algorithm) through coupled inner and outer optimizations, i.e., the result of the outer optimization model is used to measure the quality of the inner. DL-CAT devises a ground-truth construction strategy, and a pairwise loss function, allowing these two models to be trained independently; Feng et al. (2023) introduces a constrained version of BOBCAT to address the question exposure and test overlap issues. Yu et al. (2023) recently introduce the collaborative information of examinees in optimizing this bi-level problem, achieving fast convergence of proficiency estimation. Meta-Learning vs Reinforcement Learning. In CAT, meta-learning can be seen as a higher-level learning process that learns how to adapt a general strategy for question selection to specific examinees based on their responses. Actually, this meta-learning CAT can be reframed as a RL problem. Zhuang et al. (2022b) propose NCAT which transforms the meta-learning problem in CAT into a RL problem. Because the test may stop at any step according to different stopping rules, NCAT simplifies the original objective (Eq(11)) in meta-learning and sums all the test steps to minimize the loss:
(14) $\min_{\pi}\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}{\sum_{(q,y)\in\mathcal{D}_{u}^{i}}{l(y,f(q,\hat{\theta}_{i}^{t}))}}$$\triangleq$ $\mathop{\max}\limits_{\pi}\mathbb{E}_{i\sim\pi}\left[\sum_{t=1}^{T}{-{\sum_{(q,y)\in\mathcal{D}_{u}^{i}}{l(y,f(q,\hat{\theta}_{i}^{t}))}}}\right]$
$=$ $\max_{\pi}\mathbb{E}_{i\sim\pi}\left[\sum_{t=1}^{T}{-L({\hat{\theta}_{i}^{t}};\mathcal{D}_{u}^{i})}\right],$
where $\hat{\theta}_{i}^{t}=\mathop{\arg\min}_{\theta_{i}}{\sum_{(q,y)\in\mathcal{D}_{s}^{i(t)}}{l\left(y,f\left(q,\theta_{i}\right)\right)}}$ and $\mathcal{D}_{s}^{i(t)}=\{q_{1},y_{i(1)},...,q_{t},y_{i(t)}\}$. Thus, the bi-level optimization is transformed into maximizing the expected cumulative reward (i.e., $-L({\hat{\theta}_{i}^{t}};\mathcal{D}_{u}^{i})$) in RL settings, where the reward is the negative loss of the estimated proficiency of examinee $i$ on the query set at step $t$. Other RL elements (e.g., state) in NCAT are the same in Section 5.3. Recently, Wang et al. (2023a) further design a Multi-Objective RL framework of CAT. It contains three objectives in the reward function, which can improve the prediction accuracy, increase the concept diversity and reduce the question exposure respectively.
Discussion 4.
The aforementioned data-driven machine learning approaches, i.e., Reinforcement Learning and Meta Learning, are capable of directly learning from data to optimize question selection algorithms (policy). By training a policy $\pi$ on extensive response datasets and selecting $T$ questions, they strive towards achieving the ultimate objective of CAT: $\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|$. They can analyze learning behaviors, performance responses, and preferences to reveal underlying patterns and correlations. However, the training overhead for these methods is substantial, and there is also the potential risk of introducing biases present in the dataset.
5.5. Subset Selection Algorithms
The ultimate objective of CAT is to measure examinees’ abilities both efficiently and accurately. Specifically, as illustrated in Definition 1, the goal is to find a subset $S$ of $T$ questions from question bank $\mathcal{Q}$, so that the final proficiency estimate $\hat{\theta}^{T}$ can approach the true proficiency $\theta_{0}$:
(15) $\mathop{\mathrm{min}}\limits_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|.$
where $\hat{\theta}^{T}=\arg\min_{\theta}{\sum_{(q,y)\in S}{l(y,f(q,\theta))}}$ is the final proficiency estimate when the test ends $(t=T)$ with the corresponding $T$ responses. In contrast to previous sequential selection methods, it essentially does not require perfect selection at each step, but rather emphasizes the accuracy of the final estimate. From a global perspective, CAT essentially is a Subset Selection problem (Miller, 2002), a fundamental challenge in machine learning and optimization. It revolves around choosing a subset of elements $S$ from a larger set $\mathcal{Q}$ that optimizes a particular objective function $F(S)$ while adhering to specific constraints. However, we cannot directly solve the above optimization problem due to the following main challenge: The true proficiency of the examinee, denoted by $\theta_{0}$, is unknown. It is not available in the dataset, which prevents us from directly optimizing or designing the question selection algorithm. To address this, some researchers have tried to approximate the original objective. Mujtaba and Mahapatra (2021) use the standard error of measurement (SEM) to approach the optimization objective. The SEM provides a measure of confidence in an estimate from a test. They propose using SEM as a new target to measure the accuracy of proficiency estimation: $\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|\Rightarrow\min_{|S|=T}\sqrt{\frac{1}{\sum_{i\in S}{\mathcal{I}_{i}(\hat{\theta}^{T})}}},$ where $\mathcal{I}_{i}(\hat{\theta})$ is the information (e.g., Fisher Information) for a question $i$ for an examinee with proficiency $\hat{\theta}$, thus the objective function $F(S)=\sqrt{\frac{1}{\sum_{i\in S}{\mathcal{I}_{i}(\hat{\theta}^{T})}}}$. At each step, it uses multi-objective evolutionary algorithms to obtain the set of Pareto-optimal solutions (Deb, 2011) by maximizing precision and minimizing the number of questions.
Figure 8. Illustration of the subset optimization problem, adapted from (Zhuang et al., 2023): Selecting subset $S$ to cover the bank $Q$. Rectangles represent different questions, with $w(i,j)$ measuring the similarity of question pair.
To achieve a more general and scalable framework, Zhuang et al. (2023) propose a bounded estimation CAT framework (BECAT), which redefines solve the selection problem in a data summary manner. Due to the unknown $\theta_{0}$, they first find its theoretical approximation as the new target for optimization: The examinee’s one proficiency estimate $\theta^{*}$, estimated by their full responses to the entire question bank $\mathcal{Q}$, is a theoretical approximation of the true proficiency, i.e., $\theta^{*}\approx\theta_{0}$. This finding allows the selection algorithm to aim to approach $\theta^{*}$ instead of the unknown $\theta_{0}$: Select a subset of questions $S\subseteq\mathcal{Q}$, so that the examinee’s proficiency is estimated only on the subset $S$ while still (approximately) converging to the $\theta^{*}$ (i.e., the estimate that would be obtained if optimizing on the full responses to $\mathcal{Q}$).
$\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|\Rightarrow\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta^{*}\|$ $\Rightarrow\min_{|S|=T}\max_{\theta\in\Theta}\left\|\sum_{(q,y)\in S}{\gamma\nabla l(y,f(q,\theta))}-\sum_{(q,y)\in\mathcal{Q}}{\nabla l(y,f(q,\theta))}\right\|$
$\Rightarrow\mathop{\mathrm{min}}\limits_{|S|=T}\mathop{\mathrm{max}}\limits_{\theta\in\Theta}\sum_{i\in\mathcal{Q}}\mathrm{min}_{j\in S}\|\nabla l_{i}(\theta)-\nabla l_{j}(\theta)\|$
(16) $\Rightarrow\mathop{\mathrm{max}}\limits_{|S|=T}\sum_{i\in\mathcal{Q}}\mathrm{max}_{j\in S}\;w(i,j),$
where $w(i,j)\triangleq d-\mathrm{max}_{\theta\in\Theta}{\|\nabla l_{i}(\theta)-\nabla l_{j}(\theta)\|}$ is the gradient similarity between question pair $(q_{i},q_{j})$ for this examinee, thus the objective function $F(S)=\sum_{i\in\mathcal{Q}}\mathrm{max}_{j\in S}\;w(i,j)$. The core of BECAT’s subset selection strategy is to find a subset $S$ of size $T$ that maximizes the coverage of $\mathcal{Q}$, quantified by the similarity measure $w(i,j)$ between each question $i\in\mathcal{Q}$ and its closest counterpart $j\in S$. This approach, represented in Figure 8, essentially seeks the most representative questions, aligning with prior selection algorithms but under a new, more rigorous theoretical framework. Given the NP-Hard nature of this optimization, BECAT employs a submodular function approximation. A simple greedy algorithm can finally solve this subset selection problem, with BECAT ensuring that the estimate error remains upper-bounded at each step. The subset selection problem in CAT is a fresh direction with significant potential. This method offers a universal and global framework for question selection, applicable across various complex CDMs that can utilize gradient-based estimations, including neural network models. Crucially, it also provides a theoretical basis for estimation error analysis.
Table 1. Comparison of Different Question Selection Algorithms in CAT
Category Generality Interpretability Need Training Advantages Disadvantages
Statistical    Algorithms
Simple implementation and efficient operation
Dependent on IRTs and requires expert knowledge for design
Active Learning
Model-agnostic and flexible
Neglect the nuanced information within CDM parameters
Reinforcement Learning
Automatic generation of selection algorithm; Sequential Decision Making
Incurs additional training costs and potential bias from data-driven selection
Meta Learning
Automatic generation of selection algorithm; Fast Adaptation
Incurs additional training costs and potential bias from data-driven selection
Subset Selection
Strong theoretical guarantees for estimation accuracy
Faces challenges in the initial stages of CAT
Discussion 5.
Table 1 compares these five categories of selection algorithms in CAT systems, highlighting the generality and interpretability of each category, along with their main advantages and limitations. This overview assists researchers in identifying the most suitable algorithm for their CAT applications, balancing efficiency and complexity (a summary of the performance of different methods is presented in Table 3). It is noteworthy that, despite the superior performance demonstrated by the latest machine learning and deep learning methods, they do not currently supplant traditional statistical approaches. Particularly in testing scenarios that prioritize interpretability or efficiency, statistical methods remain predominant.
6. Question Bank Construction
To develop a high-quality CAT, the foundational step is to construct a high-quality question bank with a substantial number of questions. The construction of a question bank can be decomposed into two main stages as shown in Figure 9: Question Characteristics Analysis and Question Bank Development: (1) Question Characteristics Analysis first detailedly examines the properties and attributes of potential questions. Then, (2) Question Bank Development assembles the final question bank $\mathcal{Q}$ from the analyzed questions.
6.1. Question Characteristics Analysis
The first stage, question characteristics analysis, involves a detailed examination of the properties and attributes of potential questions, e.g., difficulty, discrimination, and the knowledge concepts required to answer the question (i.e., Q-matrix). To illustrate, when selecting questions based on Fisher Information, one must leverage pre-calibrated parameters like difficulty ($\beta_{j}$), discrimination ($\alpha_{j}$), and guessing factor ($c_{j}$), alongside the current proficiency estimate, to compute the Information value $\mathcal{I}_{j}(\theta)$ for each question $j$. The methods of characteristics analysis can be categorized into three main approaches: expert-based, statistic-based, and deep learning-based methods.
6.1.1. Expert-based Characteristics Annotation
In expert-based annotation, domain experts assess question parameters, a practice seen in online CAT systems like SIETTE (Conejo et al., 2004) and GenTAI (López-Cuadrado et al., 2006). Effective expert estimation often involves structured questionnaires (López-Cuadrado et al., 2008) that clearly outline the task and instructions. Follow-up discussions may refine these assessments, especially when expert opinions diverge. Results are aggregated, using averages for continuous attributes or voting for discrete ones, and sometimes applying more sophisticated statistical methods like the Consensus Theory algorithm (Kozierkiewicz-Hetmańska and Poniatowski, 2014) for expert consensus. While this approach is reliable for estimating difficulty, it can be less so for other parameters like discrimination and guessing factor (Lopez-Cuadrado et al., 2009). Furthermore, expert judgments can sometimes be subjective, potentially leading to inaccuracies in annotations, particularly when expert numbers are limited or their views are inconsistent.
6.1.2. Statistic-based Characteristics Annotation
The statistic-based method for annotating question characteristics requires gathering responses from a large group of examinees. Despite its resource-intensive nature, involving pre-testing with examinees, it utilizes parameter estimation methods or gradient descent on CDM (e.g., IRT) to estimate question parameters (De Champlain, 2010). In some psychological theories, like Classic Test Theory (CTT), question difficulty is calculated as the proportion of correct responses within examinees (Magno, 2009; DeVellis, 2006), while discrimination is derived from performance disparities between higher and lower ability examinees (Chang and Yang, 2009). The Q-matrix is another crucial characteristic of questions. It is a binary matrix that indicates which knowledge concepts are associated with a question. Numerous researchers have attempted to employ some data-driven approaches to learn the Q-matrix (Liu et al., 2013; Chen et al., 2015b). Data-driven methods like Boolean Matrix Factorization (BMF) (Sun et al., 2014) have been applied to infer it from examinee responses. The examinee response matrix ${R}$ is approximated by the element-wise multiplication of the logical complements of the examinee proficiency matrix ${A}$ and the transpose of the Q-matrix ${Q}$:
(17) ${R}=\overline{\overline{{A}}\odot{Q}^{T}}.$
The Q-matrix can then be learned by solving this equation, given the examinee response matrix and an initial estimate of the examinee proficiency matrix. An alternate recursive algorithm with a Heaviside step function is used to address the NP-hard nature of BMF and the non-invertibility of Boolean operations. Further enhancements to this method include additional judgment and processing steps to ensure convergence and robustness, as well as refinement procedures to avoid local optima in Q-matrix estimation (Xiong et al., 2022).
Figure 9. For high-quality question bank construction, two main stages are needed: question characteristics analysis and question bank development.
6.1.3. Deep Learning-based Characteristics Annotation
With the rise of Natural Language Processing (NLP), there has been an increasing trend in recent years to directly use the textual information of questions to analyze various attributes. For difficulty prediction, attention-based CNN models and domain adaptation strategies have been used to evaluate reading questions and medical question complexity (Huang et al., 2017; Qiu et al., 2019; Huang et al., 2021). For knowledge concept (Q-matrix) prediction, which typically exhibits a hierarchical structure, a Hierarchical attention-based Recurrent Neural Network has been proposed (Huang et al., 2019, 2022). Lei et al. (2021) further take into account the multi-modal features of questions, such as images and formulas. Pre-trained NLP models have also proven effective for automated question analysis (Yin et al., 2019; Ning et al., 2023).
6.2. Question Bank Development
The second stage, question bank development, involves the actual assembly of the question bank from the analyzed questions of the first stage. This process should aim to create a balanced and varied bank that can cater to different levels of proficiency and different areas of knowledge (Revuelta and Ponsoda, 1998; Reckase, 2010; Segall, 2005). This stage is critical as the quality of the question bank directly impacts the CAT system’s effectiveness. According to different usage scenarios, the approaches to developing a question bank can be categorized into the following three aspects.
6.2.1. Question Bank Blueprint Design
The goal of designing a blueprint is to generate an optimal framework for a question bank. This blueprint outlines the distribution of questions based on various statistical and non-statistical attributes associated with the questions. Reckase (2010) initially conduct a theoretical analysis of the characteristics of an optimal question bank. In a CAT system using a 1PL-IRT model with a maximum Fisher information selection algorithm, an optimal question bank is characterized by questions whose difficulty level exactly matches the estimated proficiency for each selection. However, achieving an exact match is often unfeasible. Thus, they propose a method known as the bin-and-union, which allows for a maximum deviation $r$ between optimal difficulty and estimated proficiency, is used to derive the question bank size and $r$-optimal distribution. Further research has extended these methods for large-scale CAT systems (He and Reckase, 2014) and continuous new question pretesting (Van Der Linden et al., 2000).
6.2.2. Question Bank Assembly
While the design of a question bank blueprint assumes the absence of actual questions and focuses on creating an optimal framework, the bank assembly is a process that involves generating question banks from an existing master bank according to specific requirements. Way et al. (2005) delve into the issues related to the development and maintenance of a master bank, including the constraints necessary to ensure the assembled question bank meets desired specifications. van der Linden et al. (2006) propose a mixed-integer programming model to create a question bank that satisfies all content specifications and maximizes information at selected ability values. This method, illustrated using the Law School Admission Test’s master bank, focuses on distributing estimates for the examinee population. Belov and Armstrong (2009) further introduces two significant problems related to question bank design: the direct problem and the inverse problem. The direct problem involves creating a bank that meets specific measurement requirements, such as content coverage and measurement precision, while minimizing question exposure. Conversely, the inverse problem involves determining the minimum number of questions needed to achieve the above specific requirements. A Monte Carlo method is introduced to balance measurement accuracy and question exposure.
6.2.3. Question Bank Rotating
Rotating the question bank refers to dividing a master bank into a set of smaller question banks, which may contain overlapping elements. This technique allows for the rotation of these smaller question banks among examinees, thereby ensuring a balanced exposure rate of test questions (Stocking and Swanson, 1998; Ariel et al., 2004; Swanson and Stocking, 1993). Ariel et al. (2004) proposes a method of dividing a master bank into smaller banks based on Gulliksen’s matched random subtests method (Gulliksen, 2013) to prevent issues of over- or underexposure in adaptive testing. The Weighted Deviation Model (WDM) (Swanson and Stocking, 1993) is employed to manage the degree of overlap between these banks, maintaining their representativeness while preventing question overexposure. This balance is critical to preserve the integrity of the rotation system and ensure consistent measurement across different test versions.
Discussion 6.
Think of the entire bank development process as creating and managing a library. The blueprint design (Section 6.2.1) is like the architectural plan for the library, defining where each section (e.g., fiction, non-fiction) will be located. The assembly process (Section 6.2.2) resembles acquiring books from suppliers based on specific demands. Rotating the question bank (Section 6.2.3) is similar to periodically rotating the books on display. Even though the library has a vast collection, only a subset is displayed prominently at any given time. This rotation ensures that different books get exposure, and library visitors encounter a variety of books over time. Similarly, rotating the question bank in CAT ensures a balanced exposure of test questions, enhancing the fairness and robustness of the testing process.
The construction of a high-quality question bank introduced in this section is not just a prerequisite for CAT, but also a continuous process. It requires regular updates and refinements to ensure the relevance and effectiveness of the adaptive testing system.
7. Test Control of CAT
When implementing a CAT system, in addition to considering the three components mentioned above, several key factors need to be taken into consideration. These factors include content control, fairness, robustness, and search efficiency.
7.1. Content Control
Content and exposure control is a secondary, but crucial, goal in CAT. It aims to balance the frequency of each question’s use from the question bank. Proper content control can help mitigate the risk of overexposure of questions, minimize question waste, and maximize test coverage. Two popular strategies for exposure control are the Sympson-Hetter method (Sympson and Hetter, 1985) and the A-Stratified method (Chang and Ying, 1999): (1) Sympson-Hetter Method manages question exposure rates using conditional probabilities. It doesn’t assign a selected question to the examinee immediately; instead, it passes through a probability filter. The actual chance a question is given to an examinee depends on both its selection likelihood and a content control parameter. Adjustments to the parameter aim to keep question exposure within acceptable limits. However, this method may not effectively increase the usage rate of low-exposure questions and doesn’t consider the match between question difficulty and examinee proficiency. Enhancements to this method have been developed to address these limitations (van der Linden and Veldkamp, 2004, 2007; Barrada et al., 2009b); (2) A-Stratified Method and its subsequent researches (Chang et al., 2001; Barrada et al., 2006, 2014) are designed to counteract selection biases of algorithms favoring certain questions (e.g., Fisher Information prefers highly differentiated questions). It first stratifies the question bank by difficulty, and starts with less discriminating sub-banks and progresses to more challenging ones as the test unfolds. This stratification ensures a more even exposure of all questions. On the other hand, numerous studies (Bi et al., 2020; Wang et al., 2023a) have attempted to incorporate the coverage of knowledge concepts as a criterion in question selection, aiming to make the assessment more comprehensive.
7.2. Fairness
Fairness is a topic of profound societal significance in both education and machine learning research fields, sparking numerous discussions and leading to the development of many fairness-aware learning algorithms (Cleary, 1968; Chai and Wang, 2022; Li and Liu, 2022). As a technology with potential applications in high-stakes testing, fairness in CAT is a paramount concern. In CAT, the bias that leads to fairness issues can be introduced through three components: the Cognitive Diagnosis Model (CDM), the selection algorithm, and the question bank:
  • Bias in CDMs. Biases in CDMs may stem from the skewed training data, which could reflect the underrepresentation of certain groups or pre-existing educational disparities (Liu et al., 2022; Thompson, [n.d.]; Kizilcec and Lee, 2022). Such biases can lead to an inaccurate and biased estimation of an examinee’s proficiency $\theta$, resulting in unfair outcomes.
  • Bias in Question Bank. The question bank may contain biases if questions are not equally applicable or relatable to all examinees, potentially disadvantaging certain groups (Camilli and Shepard, 1994; Hambleton et al., 1991). For example, some questions in NAPLAN have been deemed unfair for rural examinees, as these questions don’t relate to their real-life experiences (Roberts, [n.d.]). Various methods have been proposed to detect this type of bias (Chu and Lai, 2013; Camilli and Shepard, 1994; Mellenbergh, 1989).
  • Bias in Selection Algorithms. Selection algorithm can introduce bias since every algorithm has its own “selection preferences”. For example, the Maximum Fisher Information tends to select questions with high discrimination (Lord, 2012). If such questions unexpectedly correlate with specialized knowledge known only to a specific group, bias may ensue.
Figure 10. Example of a test-taker’s GRE score report. The GRE General Test provides three scaled scores: Verbal Reasoning, Quantitative Reasoning, and Analytical Writing. The percentile scores facilitate comparison with the performance of other test-takers within a specified comparison group.
Concerns about fairness in CAT also stem from the fact that examinees answer different questions (Green et al., 1984). Equating, a technique used to ensure score equivalence across different tests, is commonly employed to address such concerns (Brigman and Bashaw, 1976). Many further studies about equating scores have been conducted (van der Linden, 2000; JAN, [n.d.]; Sawaki, 2001). Equating ensures fairness by satisfying three requirements (van der Linden, 2000; Braun, 1982): (1) Measurement of the Same Variable: the proficiency estimated by the CAT must reflect the same aspects of proficiency; (2) Equity: an examinee’s score on one test after equating should match the score they would have received if they had taken this test directly; (3) Invariance across Populations: every examinee should get the same equated score for the same performance in different populations. In real-world tests such as the GRE, equating has been used to standardize scores and percentiles (Figure 10), taking into account the difficulty of the questions answered. This process ensures that scores can be compared fairly across different examinees worldwide.
7.3. Robustness
Noise in CAT can impact the precision of the estimated proficiency of an examinee, leading to potential errors in score interpretation. In CAT, noise usually refers to the random variability or measurement error that can affect the accuracy of estimation. It can arise from various sources such as test administration conditions, test-taker behavior, or question characteristics. To mitigate the effects of noise in CAT, a robustness factor is introduced to help stabilize the estimation of proficiency by incorporating additional information, thereby counteracting the impact of noise and improving the reliability (Baker and Kim, 2004). In machine learning, various robustness techniques are employed to enhance the performance of models in the presence of noise, such as regularization methods (Li and Zhang, 2021), data augmentation (Wang et al., 2021), adversarial methods (Dong et al., 2020), ensemble methods (Kariyappa and Qureshi, 2019). In the CAT testing process, significant sources of noise such as guess and slip factors made by examinees, introduce uncertainty. For example, an examinee’s proficiency level may not be uniquely determined by their responses, as they may solve a particular question correctly using different knowledge concepts or even by guessing. The presence of this noise and uncertainty poses a significant challenge to the robustness of CAT. Veldkamp and Verschoor (2019) consider the uncertainty in the question parameters in the selection process. Recently, Zhuang et al. (2022a) explores the use of ensemble learning to combine multiple potential estimates at each step to enhance the proficiency estimation.
7.4. Search Efficiency
In large-scale educational testing, such as the GRE (Mills and Steffen, 2000), the search efficiency of selection algorithms poses a significant challenge. The algorithms typically evaluate the value of each question in a brute-force manner. Consequently, the time complexity of selecting a question in each testing step is linear to the number of questions in the bank $\mathcal{Q}$, denoted as $O(|\mathcal{Q}|)$. This linear relationship is a bottleneck in intelligent educational systems. Many organizations like GMAT (Rudner, 2009) resort to the manual development of filtering rules by experts, a labor-intensive process that somewhat limits the efficacy of CAT. To address the efficiency issue in question selection, researchers have proposed two kinds of solutions: (1) The application of the Particle Swarm Optimization (PSO) algorithm in IRT has proven effective in mitigating computational complexities (Huang et al., 2009). PSO operates with particles concurrently exploring the search space, where each particle represents a potential solution to the optimization problem of question selection. This parallel exploration allows PSO to cover a larger portion of the search space in each iteration, improving the chances of rapidly identifying optimal questions; (2) Leveraging efficient search methods from the realms of recommendation systems and information retrieval to enhance CAT’s search efficiency, e.g., creating of a tree-based index (Ram and Gray, 2012; Zhu et al., 2018). This approach organizes data in hierarchical tree structures, with balanced trees ensuring logarithmic time complexity for search operations. Recently, Hong et al. (2023) propose a Search-Efficient CAT framework, which employs examinee-aware space partitioning to construct a tree-based index. It reduces the search space and eliminates redundant searches in multi-round testing steps, thereby reducing the search complexity from $O(|\mathcal{Q}|)$ to $O(\log|\mathcal{Q}|)$.
Discussion 7.
Table 2 showcases the underlying causes and advantages of different factors in CAT test control. While accuracy and efficiency stand as primary objectives, these factors hold significant importance for practical settings, especially in high-stakes testing scenarios (e.g., competitive or selective examinations). However, consideration of these factors may inevitably reduce accuracy. For example, when considering the additional fairness to ensure equity among different groups, it might be necessary to deviate from the optimal trajectory of a well-trained selection algorithm. Thus, CAT poses a multidimensional decision-making challenge, necessitating the consideration of various factors at the same time using diverse machine learning techniques.
Table 2. Test Control: Key Factors in CAT Implementation
Factors Category Causes Advantages Pubs
Content Control
Exposure
Control
Unbalanced question usage
Mitigates overexposure;
Test security
(Sympson and Hetter, 1985; Chang and Ying, 1999)
(Barrada et al., 2009b, 2014)
Test Coverage Algorithmic preferences
Comprehensive assessment;
Broad knowledge coverage
(Bi et al., 2020; Wang et al., 2023a)
Fairness Bias in CDMs
Skewed training data;
Underrepresentation of certain groups
Promotes equitable outcomes;
Improves accuracy of proficiency estimation
(Kizilcec and Lee, 2022; Thompson, [n.d.])
(Liu et al., 2022)
Bias in Question
Bank
Unequal applicability;
Cultural or regional biases
Ensures content relevance;
Reduces disadvantage for certain groups
(Camilli and Shepard, 1994; Hambleton et al., 1991)
(Roberts, [n.d.]; Chu and Lai, 2013)
Bias in Selection
Algorithms
Algorithmic preferences
Reduces disadvantage for certain groups
(Lord, 2012)
Equating
Different selected questions across examinees
Score comparability;
Fairness across different tests
(Green et al., 1984; van der Linden, 2000)
(JAN, [n.d.]; Sawaki, 2001)
Robustness Noise Resistance
Random variability;
Guessing and slipping factors
Stabilizes estimation;
Improves reliability
(Baker and Kim, 2004; Zhuang et al., 2022a)
Modeling Uncertainty
Uncertainty in response
Improves accuracy of proficiency estimation
(Veldkamp and Verschoor, 2019)
(Zhuang et al., 2022a)
Search Efficiency
Large question banks;
Brute-force search
Reduces search complexity (Huang et al., 2009; Hong et al., 2023)
8. Evaluation
CAT aims to estimate examinees’ proficiencies with high accuracy and efficiency. While the true proficiency is inherently unobservable and cannot serve as a direct benchmark, various metrics have been developed to assess the performance of CAT methods, such as correlation coefficients, bias, and measurement error (Crocker and Algina, 1986; Van der Linden and Glas, 2010). This section introduces two of the most extensively utilized evaluation methods: simulation of proficiency estimation and examinee score prediction.
Simulation of Proficiency Estimation.
The simulation of ability estimation is a foundational evaluation technique in CAT (Vie et al., 2017). Since we don’t have direct access to an examinee’s true proficiency level $\theta_{0}$, we create a simulated version of it. By sampling a set of true proficiency values $\{\theta_{0}^{1},\theta_{0}^{2},...,\theta_{0}^{N}\}$ from the distribution of examinee proficiencies, we construct a virtual cohort of $N$ ”examinees”. This approach enables us to further emulate the interactions between examinees (with these proficiencies) and any question from the question bank, utilizing CDMs. Consequently, upon the conclusion of a CAT assessment, the estimated proficiency values $\hat{\theta}^{T}$ can be directly compared with the true values $\theta_{0}$. For instance, by computing the Mean Square Error (MSE) between these sets of values $\mathbb{E}\|\hat{\theta}^{t}-\theta_{0}\|$, we gain an accurate assessment of the CAT system’s precision in estimating examinee proficiencies (Bi et al., 2020; Cheng, 2009).
Examinee Score Prediction.
In CAT systems powered by machine learning, proficiency estimates are validated by predicting examinee responses (correct or incorrect) on a held-out test set. For example, we can allocate 70%-20%-10% examinees for training, validation, and testing respectively, and the examinees in the validation/testing set won’t appear in training. Training data calibrates question parameters (Sections 6.1) and optimizes data-driven selection algorithms (Sections 5.3, 5.4). During validation or testing, the responses of each examinee are further divided into a candidate set $\mathcal{Q}_{i}$ for selecting questions and a held-out meta set $\mathcal{M}_{i}$ for evaluation. The candidate set $\mathcal{Q}_{i}$ (with corresponding response label $y$) is used to simulate the CAT procedure: It involves selecting questions from $\mathcal{Q}_{i}$, updating proficiency estimates after each step, and then accessing estimate’s precision by predicting responses on $\mathcal{M}_{i}$. The assumption is that better score predictions reflect more accurate proficiency estimates. Thus, the binary classification metrics can be used for evaluations, e.g., Prediction Accuracy (ACC) (Gao et al., 2021) and Area Under ROC Curve (AUC) (Bradley, 1997).
8.1. Datasets
To rigorously assess the efficacy of a CAT solution, it is essential to deploy diverse datasets that can effectively challenge and evaluate the system’s capabilities. Three types of data can be employed for this evaluation: (1) Educational Assessment Datasets: These datasets originate from actual standardized tests such as the SAT, GRE, or GMAT. They encompass a broad spectrum of questions covering multiple subjects, coupled with responses from a substantial pool of test-takers. Such datasets are invaluable for gauging the CAT system’s capacity. However, due to stringent privacy laws, intellectual property rights, and the protective measures taken by educational testing companies regarding their questions and datasets, most of these are proprietary and not easily accessible. (2) Real-World Educational Data: This category includes data collected from educational environments in practice, such as schools, universities, and online learning platforms. It provides insights into how examinees interact with educational content and assessments in a natural setting. The data may encompass examinee information, performance responses, learning behaviors, question characteristics, etc. We have open-sourced a comprehensive education-related dataset library: https://github.com/bigdata-ustc/EduData. It includes a range of publicly available datasets along with previously private datasets. Additionally, we have provided a detailed data analysis to support further CAT research and application in educational settings. The following are introductions to several commonly used datasets, and more datasets can be found at the EduData gitHub link.
  • ASSISTments (Feng et al., 2009), established in 2004, is an online tutoring platform in the United States that offers examinees both assessments and instructional support. To date, the ASSISTments team has released four public datasets [^2] https://sites.google.com/site/assistmentsdata/datasets/: ASSISTments2009, ASSISTments2012, ASSISTments2015, and ASSISTments2017. These datasets are response data and mostly collected from mathematics in middle school. They also include valuable side information, such as attempt count (the number of tries an examinee has made), ms first response (the time it takes for an examinee’s first response), problem type, and average confidence.
  • Junyi Dataset (Chang et al., 2015) includes logs and exercise data from Junyi Academy, a Chinese online learning platform launched in 2012 using Khan Academy’s open-source code. It features a detailed question hierarchy and relationships, labeled by experts.
  • MOOCCube [^3] https://www.biendata.xyz/competition/chaindream_mooccube_task2/, Massive Open Online Courses (MOOCs) are among the most prevalent platforms for online learning. This dataset collects examinees’ responses to questions related to various computer science knowledge concepts. Additionally, the dataset includes the text of the problems, which can be used to enhance the performance of question selection, proficiency estimation, question characteristics analysis, etc.
  • EdNet Dataset (Choi et al., 2020) is a large collection of examinee learning records from the AI tutoring system Santa [^4] https://github.com/riiid/ednet, which is used for English language learning in South Korea. It focuses on examinees preparing for the eTOEIC (Test of English for International Communication) Listening and Reading Test, with over 131 million learning records from approximately 784,000 examinees.
  • Eedi2020 Dataset (Wang et al., 2020a), released for the NeurIPS 2020 Education Challenge, contains over 17 million records of examinees’ responses to mathematics multiple-choice questions on the Eedi platform [^5] https://eedi.com/projects/neurips-education-challenge. It includes detailed information on examinees’ choices, demographics, and containment relationships of knowledge concepts, as well as associated quiz and curriculum metadata. This extensive dataset enables in-depth analysis of examinee behaviors and the development of personalized tools.
(3) Simulated Datasets. These are artificially created datasets that mimic the characteristics of real examinee responses as illustrated above. They can be tailored to include specific patterns, noise levels, and distributions, allowing for controlled testing of the CAT system under various scenarios. Monte Carlo simulations can also be used to generate datasets with known properties and ground truth (Van der Linden and Glas, 2010). These datasets are useful for validating the CAT system’s capability to estimate proficiencies accurately and to adapt to the simulated changes during the testing process.
Table 3. AUC results reported by different CAT methods
ASSISTments Eedi2020
Selection Algorithm CDM IRT (Embretson and Reise, 2013) NeuralCD (Wang et al., 2020c) IRT (Embretson and Reise, 2013) NeuralCD (Wang et al., 2020c)
AUC@5 AUC@20 AUC@5 AUC@20 AUC@5 AUC@20 AUC@5 AUC@20
Random
67.68 (Wang et al., 2023a)
70.68 (Zhuang et al., 2023)
65.86 (Ma et al., 2023)
68.43 (Wang et al., 2023a)
72.61 (Zhuang et al., 2023)
67.73 (Wang et al., 2023a)
71.19 (Zhuang et al., 2023)
70.52 (Yu et al., 2023)
69.70 (Wang et al., 2023a)
72.83 (Zhuang et al., 2023)
68.38 (Wang et al., 2023a)
69.05 (Zhuang et al., 2023)
71.98 (Wang et al., 2023a)
74.82 (Zhuang et al., 2023)
68.45 (Wang et al., 2023a)
69.32 (Zhuang et al., 2023)
73.67 (Yu et al., 2023)
72.98 (Wang et al., 2023a)
74.99 (Zhuang et al., 2023)
Statistical Algorithms Fisher Information (Lord, 2012)
67.95 (Wang et al., 2023a)
71.33 (Zhuang et al., 2023)
66.41 (Ma et al., 2023)
69.26 (Wang et al., 2023a)
73.54 (Zhuang et al., 2023)
68.92 (Wang et al., 2023a)
70.60 (Zhuang et al., 2023)
72.66 (Wang et al., 2023a)
76.24 (Zhuang et al., 2023)
KL Information (Chang and Ying, 1996)
67.92 (Wang et al., 2023a)
71.38 (Zhuang et al., 2023)
69.23 (Wang et al., 2023a)
73.57 (Zhuang et al., 2023)
68.69 (Wang et al., 2023a)
69.79 (Zhuang et al., 2023)
72.60 (Wang et al., 2023a)
75.73 (Zhuang et al., 2023)
Fisher Information
+ Robust (Zhuang et al., 2022a)
68.93 (Zhuang et al., 2022a) 75.99 (Zhuang et al., 2022a)
KL Information
+ Robust (Zhuang et al., 2022a)
68.90 (Zhuang et al., 2022a) 76.03 (Zhuang et al., 2022a)
Active Learning MAAT (Bi et al., 2020)
68.24 (Wang et al., 2023a)
71.54 (Zhuang et al., 2023)
66.24 (Ma et al., 2023)
69.7 (Wang et al., 2023a)
73.08 (Zhuang et al., 2023)
67.96 (Wang et al., 2023a)
70.98 (Zhuang et al., 2023)
70.85 (Yu et al., 2023)
71.17 (Wang et al., 2023a)
72.27 (Zhuang et al., 2023)
69.09 (Wang et al., 2023a)
70.32 (Zhuang et al., 2023)
73.19 (Wang et al., 2023a)
74.46 (Zhuang et al., 2023)
69.03 (Wang et al., 2023a)
70.12 (Zhuang et al., 2023)
74.33 (Yu et al., 2023)
73.75 (Wang et al., 2023a)
75.83 (Zhuang et al., 2023)
MAAT
+ Robust (Zhuang et al., 2022a)
68.93 (Zhuang et al., 2022a) 76.09 (Zhuang et al., 2022a) 70.39 (Zhuang et al., 2022a) 76.63 (Zhuang et al., 2022a)
Reinforcement Learning GMOCAT (Wang et al., 2023a) 69.13 (Wang et al., 2023a) 71.91 (Wang et al., 2023a) 69.95 (Wang et al., 2023a) 72.95 (Wang et al., 2023a) 69.81 (Wang et al., 2023a) 74.19 (Wang et al., 2023a) 71.25 (Wang et al., 2023a) 75.76 (Wang et al., 2023a)
NCAT (Zhuang et al., 2022b)
68.67 (Wang et al., 2023a)
71.53 (Zhuang et al., 2023)
71.06 (Wang et al., 2023a)
73.50 (Zhuang et al., 2023)
69.28 (Wang et al., 2023a)
71.59 (Zhuang et al., 2023)
72.53 (Yu et al., 2023)
71.68 (Wang et al., 2023a)
73.59 (Zhuang et al., 2023)
69.04 (Wang et al., 2023a)
72.11 (Zhuang et al., 2023)
73.32 (Wang et al., 2023a)
76.66 (Zhuang et al., 2023)
69.09 (Wang et al., 2023a)
74.10 (Zhuang et al., 2023)
74.49 (Yu et al., 2023)
74.55 (Wang et al., 2023a)
79.12 (Zhuang et al., 2023)
Meta Learning Algorithms BOBCAT (Ghosh and Lan, 2021)
68.65 (Wang et al., 2023a)
71.68 (Zhuang et al., 2023)
66.41 (Ma et al., 2023)
70.97 (Wang et al., 2023a)
73.39 (Zhuang et al., 2023)
69.50 (Wang et al., 2023a)
71.45 (Zhuang et al., 2023)
71.98 (Yu et al., 2023)
71.80 (Wang et al., 2023a)
72.84 (Zhuang et al., 2023)
68.94 (Wang et al., 2023a)
74.42 (Zhuang et al., 2023)
73.24 (Wang et al., 2023a)
76.58 (Zhuang et al., 2023)
69.17 (Wang et al., 2023a)
76.00 (Zhuang et al., 2023)
75.12 (Yu et al., 2023)
74.51 (Wang et al., 2023a)
79.00 (Zhuang et al., 2023)
DL-CAT (Ma et al., 2023) 66.68 (Ma et al., 2023)
SACAT (Yu et al., 2023) 75.24 (Yu et al., 2023) 75.48 (Yu et al., 2023)
Subset Selection
Algorithms
BECAT (Zhuang et al., 2023) 71.44 (Zhuang et al., 2023) 73.61 (Zhuang et al., 2023) 71.60 (Zhuang et al., 2023) 73.70 (Zhuang et al., 2023) 73.15 (Zhuang et al., 2023) 76.82 (Zhuang et al., 2023) 76.30 (Zhuang et al., 2023) 79.36 (Zhuang et al., 2023)
Table 3 displays some representative methods of each category of selection algorithms and their AUC results on two different datasets. The comparison in this survey focuses on the results at the early testing stage (step=5) and the final testing stage (step=20). It is important to note that the results cannot be directly compared if experimental settings are not standardized. Despite this, the table as a whole reveals that data-driven methods (e.g., reinforcement learning, meta-learning methods) generally outperform statistical methods. This is because these methods can train and optimize selection algorithms from examinee large-scale response data, while statistical methods simply adhere to fixed functions for selecting questions. The latest subset selection methods do not require training but are remarkably effective. This is primarily because they attempt to explicitly approximate the objectives of CAT and provide theoretical guarantees on estimation errors. Furthermore, it is observed that considering factors within test control, such as robustness, can enhance accuracy.
9. Opportunities for Future Research
The integration of machine learning into CAT is poised to revolutionize the field. This section explores the future potential of machine learning to expand the applicability, interpretability, and multi-dimensionality of CAT systems.
Multi-Dimensionality of the Assessment Process
Future research should harness machine learning to enhance the multi-dimensionality of the assessment. This involves not only the traditional response patterns but also the nuanced analysis of process data such as response times and mouse movements, which can provide insights into an examinee’s problem-solving strategies and levels of engagement. Moreover, integrating learning data, such as the examinee’s prior interactions with materials, can offer a longitudinal perspective on their learning trajectory and readiness for new concepts. Additionally, the analysis of educational content, encompassing textual, visual, and auditory materials (Xiao et al., 2022; Shi et al., 2014), allows for a richer understanding of how examinees interact with multifaceted information. Such machine learning-driven approaches promise to refine CAT systems comprehensively, enabling them to deliver assessments that are not just accurate reflections of an examinee’s proficiency but also predictive of their potential for future learning.
Multi-Stage Testing
Multi-Stage Testing (MST) is a variant of CAT that is frequently employed in high-stakes examinations (Magis et al., 2017). Unlike CAT, which select a single question at each step, MST presents a set of questions (i.e., testlet) at each step, which is pre-designed by experts as the unit of selection. The selection process is predetermined by the test design, based on a decision tree or a similar algorithmic structure, aiming to match test takers’ estimated proficiency levels more closely (Mead, 2006). Building upon CAT, MST presents more complex challenges, including considerations around the number of testlets, the size of each testlet, and questions’ combinatory effects within a testlet. Currently, the selection algorithm for MST relies heavily on manual design. To facilitate the widespread adoption of MST for large-scale testing, there is a need for increased research into the automated MST construction.
Empowering CAT with Generative AI
Generative artificial intelligence, with Large Language Models (LLMs) at the forefront, is trained on massive, cross-domain datasets, endowing it with versatility and a profound repository of world knowledge (Dong et al., 2023; Chang et al., 2023). These models have already shown preliminary progress in user modeling, such as recommendation systems, and in the generation of personalized strategy. In the future, there is potential for these large models to significantly enhance CAT systems in various aspects, such as question selection, proficiency assessment, and even the automatic generation of novel, tailored questions on the fly (Bhandari et al., 2023) – questions that are not pre-existing in the bank. This could lead to more effective and individualized testing experiences.
Towards Explainable Machine Learning in CAT
Traditional CAT systems, particularly those based on information and statistic approaches, are lauded for their interpretability, from the parameters of CDMs to the logic behind the question selection algorithm. This transparency provides valuable insights to all stakeholders, including examinees, parents, and educators, and supports developers in debugging and refining the CAT system behind the final testing scores. However, recent machine learning approaches, especially those employing deep learning, have an overwhelming advantage in capabilities on knowledge discovery, at the cost of reduced interpretability. Bridging the gap between these paradigms to create CAT systems that are both accurate and self-explanatory is a significant challenge that future research must address. This is particularly crucial for high-stakes standardized testing, where the outcomes carry significant consequences
CAT for AI System Evaluation
Computerized Adaptive Testing (CAT) has traditionally been tailored for educational purposes, primarily serving humans. However, the intrinsic qualities of CAT—its efficiency at discerning test-taker’s cognitive states and its capacity for conducting multi-step, adaptive interactions—suggest a broader potential. Given these capabilities, we can foresee its adaptation to not only serve a wider human audience but also to assess and refine AI systems. In the domain of AI, CAT methodologies could become crucial in the efficient evaluation of machine learning models, including their cognitive capabilities. This is particularly relevant in the advanced Large Language Models (LLMs) that are approaching human-level performance, where CAT could provide nuanced insights into their cognitive-like functions. Moreover, CAT could expedite model capacity evaluation using the least possible amount of samples. This evolution can deepen our understanding of both human and machine cognition, paving the way for innovative uses in efficient evaluation services, and smarter, more responsive AI systems.
10. Conclusion
This comprehensive survey underscored the intricate and vast system that is Computerized Adaptive Testing (CAT), emphasizing the potential and prospects of applying machine learning to enhance CAT. The paper focused on CAT’s two primary concerns of accuracy and efficiency within the assessment process from the machine learning perspective. The presented insights were accessible and relevant not just for specialists in education or psychometrics but for a wide array of researchers. We encourage interested readers to learn about the transformative impact of machine learning in this field and take this survey as a reference for future research.